# AWP: Activation-Aware Weight Pruning and Quantization with Projected Gradient Descent

### v1

@@ 1 Introduction
In this work, we consider Post-Training Quantization (PTQ) and pruning/sparsification. This is in contrast to Quantization-Aware Training (QAT) and sparse network training of large models which are still resource intensive, and many applications only consider leveraging pre-trained foundation models instead of training a model from scratch. In the post-training compression setup (Nagel et al., 2020; Li et al., 2021; Hubara et al., 2021; Liang et al., 2021), a trained but uncompressed model is given, together with a small amount of calibration data, with the aim to produce an accurate compressed model without retraining. As the foundation models are very large, which may exceed the hardware memory limit during compression, recent works, such as Wang et al. (2020); Nagel et al. (2020); Hubara et al. (2021), further break the compression task into layer-wise sub-problems, identifying a compressed weight approximation for each layer, given a sub-sample of the layer’s inputs and outputs based on calibration data. This approach is the focus of our work.
@@ 2 Related Work and Background
In both the post-training quantization and pruning literature, activation-aware methods, e.g., Lin et al. (2024); Frantar et al. (2022a); Wang et al. (2024a); Frantar & Alistarh (2023); Sun et al. (2023); Frantar et al. (2022b), demonstrate superior performance as they consider the input activation statistics to guide the quantization and pruning. For example, the traditional magnitude based weight pruning essentially solves a sparse approximation problem, without considering input activations:
@@
However, there is no closed-form solution to the above problem and several heuristic methods have been proposed. Wanda (Sun et al., 2023) computes the $\ell_{2}$-norm of each row of $\mathbf{X}$, i.e., $\|\mathbf{X}[i,:]\|_{2}$, for $i=1,\ldots,d_{\text{in}}$, to scale the corresponding column of $\mathbf{W}$, and then performs magnitude based pruning on this scaled matrix to obtain the pruning mask for the weight $\mathbf{W}$. It can be viewed as approximating $\mathbf{C}^{\frac{1}{2}}$ only by its diagonal in Equation 3. Therefore, the cross-correlation information of the input activations $\mathbf{X}$ between each dimension is completely discarded.
Similar to Wanda, in the LLM quantization literature, the state-of-the-art Activation-aware Weight Quantization (AWQ) method (Lin et al., 2024) computes the scaled $\ell_{1}$-norm of each row of $\mathbf{X}$, i.e., $\|\mathbf{X}[i,:]\|_{1}/n$, for $i=1,\ldots,d_{\text{in}}$, to scale the corresponding column of $\mathbf{W}$, and performs quantization on the scaled matrix.
@@
The Optimal Brain Compression (OBC) method (Frantar et al., 2022b) simplifies the original Optimal Brain Surgeon (OBS) (LeCun et al., 1989; Hassibi et al., 1993) by breaking the compression task into layer-wise sub-problems. For each layer, they use a greedy solver that sequentially prunes (or quantizes) weights. To prune each row of $\mathbf{W}$, they iteratively zero out the entry that results in the smallest incremental increase to approximation loss. However, their follow-up work (Frantar & Alistarh, 2023) mentions that this method is hard to scale to models with billions of parameters, and further propose several approximations (e.g., prune the weights from left to right and only recalculate approximation loss impact for the weights to the right) for speedup, leading to the SparseGPT method (Frantar & Alistarh, 2023) for pruning, and the GPTQ (Frantar et al., 2022a) method for quantization.
Due to space limits, we refer interested readers to Zhu et al. (2024); Gholami et al. (2022); Nagel et al. (2021); Kim et al. (2023); Wan et al. (2023) for a comprehensive overview of weight quantization and pruning methods. Another line of work on structured pruning, e.g., Ma et al. (2023); Xia et al. (2024), removes entire structured components of a network, but usually involves post-training or fine-tuning to recover the performance.
@@ 3 Motivation and Method
Many existing sparse approximation and compressive sensing algorithms can be used to solve this problem, such as Matching Pursuit (Mallat & Zhang, 1993), Orthogonal Matching Pursuit (OMP) (Cai & Wang, 2011), Iterative Hard Thresholding (IHT) (Blumensath & Davies, 2009), CoSaMP (Tropp & Needell, 2008), Basis Pursuit (Chen et al., 2001), Sparse Bayesian Learning methods (Tipping, 2001; Wipf & Rao, 2004). In fact, the OBC method can be viewed as the OMP method in reverse order. However, it is well-known in Compressive Sensing that such methods are too greedy and often outperformed by IHT and $\ell_{1}$ type methods.
We adopt the IHT method, since it is efficient to run and also has recovery guarantees (Blumensath & Davies, 2009). IHT is a special form of the projected gradient descent (PGD) method. Each iteration of PGD involves calculating the gradient [^3] Using the fact that $\mathbf{y}=(\mathbf{W}[i,:]\mathbf{C}^{\frac{1}{2}})^{\top}$ can simplify the gradient computation without doing SVD of $\mathbf{C}$, as explained later. of $f(\boldsymbol{\theta})$, i.e., $\nabla f(\boldsymbol{\theta}^{(t)})$, and projecting the updated weights $(\boldsymbol{\theta}^{(t)}-\eta\nabla f(\boldsymbol{\theta}^{(t)}))$ onto the constraint set. For IHT, the projection onto the constraint $\|\boldsymbol{\theta}\|_{0}\leq k$ is simply hard-thresholding, i.e., keeping the $k$ largest-magnitude elements of ($\boldsymbol{\theta}^{(t)}-\eta\nabla f(\boldsymbol{\theta}^{(t)})$, and setting the remaining elements to 00. Actually all the terms in (4) can be solved independently and in parallel. This can also be viewed as minimizing the following objective, with the constraint of (5) that each row of $\hat{\mathbf{W}}$ is $k$-sparse:
@@
$\nabla f_{1}(\hat{\mathbf{W}})=$ $-2(\mathbf{W}\mathbf{C}^{\frac{1}{2}}-\hat{\mathbf{W}}\mathbf{C}^{\frac{1}{2}})(\mathbf{C}^{\frac{1}{2}})^{\top}$ (8)
@@
[…]
@@ 4 Experiments
We compare the proposed AWP method with state-of-the-art methods on pruning, quantization, as well as joint pruning and quantization. As in AWQ and Wanda, we evaluate the perplexity of the compressed model on the held-out WikiText-2 (Merity et al., 2016) validation set.
@@ 4.1 Pruning
To compare with Wanda, SparseGPT, and Magnitude-based pruning, we follow the exact setup of the experiments in Wanda [^4] https://github.com/locuslab/wanda. More specifically, we tested on the Llama-2 7B and 13B models. The calibration data $\mathbf{X}$ is 128 sequences (each has 2048 tokens) sampled from the C4 training set (Raffel et al., 2020). We test pruning ratios from $\{50\%,60\%,70\%,80\%,90\%\}$. Note that larger pruning ratio corresponds to higher compression rate. For AWP, the iteration stops when the Frobenius norm of the gradient normalized by the Frobenius norm of the original weight is less than 0.0001, or 200 iterations is reached. The step size $\eta$ is set as $2/\|\mathbf{C}\|_{\mathrm{F}}$. As problem (3) is nonconvex, we initialize $\mathbf{\Theta}^{(0)}$ as the solution of Wanda, since a good initial point helps nonconvex optimization.
Table 1 shows the perplexity of the pruned Llama-2-7B model for different methods and pruning ratios, and Table 2 shows the corresponding results for the Llama-2-13B model. The results of SparseGPT and Magnitude-based pruning are from Sun et al. (2023). As a reference, the perplexity of the original Llama-2-7B dense model (i.e., pruning ratio = 0%) is 5.12 and the perplexity of the original Llama-2-13B dense model is 4.57.
@@
Magnitude $14.89$ $34e$ 3 - NaN -
SparseGPT $6.51$ $9.58$ $21e$ 2 -
Wanda $6.48$ $10.09$ $70.04$ $34e$ 3 $41e$ 4
@@
Magnitude $6.37$ $11.23$ $45e$ 4 -
SparseGPT $5.63$ $7.80$ $21e$ 2 -
Wanda $5.59$ $7.97$ $43.06$ $31e$ 3 $42e$ 4
@@ 4.2 Quantization
We compare with state-of-the-art layer-wise weight quantization methods AWQ [^5] https://github.com/mit-han-lab/llm-awq and GPTQ [^6] https://github.com/AutoGPTQ/AutoGPTQ/tree/main. We follow the experiment setups of AWQ, which focus on weight-only grouped quantization with a group size of 128. As in AWQ, we use a small calibration set from the Pile dataset (Gao et al., 2020) in order not to overfit to a specific downstream domain. Besides INT4 and INT3 quantization, we additionally experiment with INT2 quantization.
@@ 4.3 Joint Pruning and Quantization
An interesting finding is that, compared with INT2 quantization in Table 3, INT4 quantization with 75% pruning ratio achieves significantly better perplexity. Note that 4 bits combined with 75% pruning rate is roughly equivalent to 2 bits, as we need 1 bit to store the pruning mask. As recent efforts try to push the quantization to extremely low bits, e.g., 1.58 bit (Ma et al., 2024; Wang et al., 2024b), our results show that combining quantization with pruning may achieve much better performance.
@@ 5 Conclusion and Discussion
Our future directions include extending pruning to structured sparsity, e.g., NVIDIA’s 2:4 sparsity (Mishra et al., 2021), though the current unstructured sparsity can already be accelerated by GPUs like Cerebras Wafer-Scale Engine [^7] https://www.cerebras.net/blog/harnessing-the-power-of-sparsity-for-large-gpt-ai-models. Another direction is to provide theoretical guarantees of the proposed method for quantization.
@@ A.1 Convergence under Restricted Isometry Property (RIP)
We first recall the theoretical guarantees of IHT algorithm (Blumensath & Davies, 2009), which uses the iteration:
@@
[Blumensath & Davies (2009)] Given a noisy observation $\mathbf{y}=\mathbf{A}\boldsymbol{\theta}_{k}+\mathbf{e}$, where $\boldsymbol{\theta}_{k}$ is $k$-sparse. If $\mathbf{A}$ has the restricted isometry property with $\beta_{3k}<1/8$, then, at iteration $t$, IHT will recover an approximation $\boldsymbol{\theta}^{(t)}$ satisfying
@@
The definition of $\beta_{3k}$ in RIP can be found in Equation (6) of Blumensath & Davies (2009).
@@ A.2 Convergence under Restricted Strong Convexity (RSC) and Restricted Smoothness (RSM) Property
On the other hand, instead of requiring the restricted isometry property (RIP) of $\mathbf{C}^{\frac{1}{2}}$, which is not easy to verify (Wang et al., 2016), Jain et al. (2014) provided the recovery guarantee of IHT method based on Restricted Strong Convexity (RSC) and Restricted Smoothness (RSM) properties defined below (Jain et al., 2014; Liu & Foygel Barber, 2019):
@@
Based on the above two properties,  Liu & Foygel Barber (2019) states that for an objective function $f$ satisfying $(\alpha,k)$-RSC and $(\beta,k)$-RSM, IHT with step size $\eta\propto 1/\beta$, i.e., ${\boldsymbol{\theta}}^{(t+1)}=H_{k}({\boldsymbol{\theta}}^{(t)}-\eta\nabla_{\boldsymbol{\theta}}f({\boldsymbol{\theta}}^{(t)}))$, satisfies
@@
Note that, in Definition A.4 and Definition A.5, $\ \|{\boldsymbol{\theta}}_{1}\|_{0}\leq k$ and $\|{\boldsymbol{\theta}}_{2}\|_{0}\leq k$, therefore $\|{\boldsymbol{\theta}}_{1}-{\boldsymbol{\theta}}_{2}\|_{0}\leq\min(2k,d_{\mathrm{in}})$, which is still a sparse vector when $k<d_{\mathrm{in}}/2$. So $2\lambda_{\min}(\mathbf{C})$ and $2\lambda_{\max}(\mathbf{C})$ are usually loose lower and upper bounds for $\alpha$ and $\beta$, respectively, and $\lambda_{\max}(\mathbf{C})/\lambda_{\min}(\mathbf{C})$ is usually a loose upper bound for $\kappa$. One can see that, the smaller the $\kappa$, the better the convergence guarantee and recovery guarantee.

### v2

@@ 1 Introduction
In this work, we consider Post-Training Quantization (PTQ) and pruning/sparsification. This is in contrast to Quantization-Aware Training (QAT) and sparse network training of large models which are still resource intensive, and many applications only consider leveraging pre-trained foundation models instead of training a model from scratch. In the post-training compression setup (nagel2020up; DBLP:journals/corr/abs-2102-05426; pmlr-v139-hubara21a; LIANG2021370), a trained but uncompressed model is given, together with a small amount of calibration data, with the aim to produce an accurate compressed model without retraining. As the foundation models are very large, which may exceed the hardware memory limit during compression, recent works, such as pmlr-v119-wang20c; nagel2020up; pmlr-v139-hubara21a, further break the compression task into layer-wise sub-problems, identifying a compressed weight approximation for each layer, given a sub-sample of the layer’s inputs and outputs based on calibration data. This approach is the focus of our work.
@@ 2 Related Work and Background
For the post-training quantization and pruning, activation-aware methods, e.g., lin2024awq; frantar2022gptq; wang2024q; frantar2023sparsegpt; sun2023wanda; 10.5555/3600270.3600593; ZHANG2025101778, demonstrate superior performance as they consider the input activation statistics to guide the quantization and pruning. For example, the traditional magnitude based weight pruning essentially solves a sparse approximation problem, without considering input activations:
@@
However, there is no closed-form solution to the above problem and several heuristic methods have been proposed. Wanda (sun2023wanda) computes the $\ell_{2}$-norm of each row of $\mathbf{X}$, i.e., $\|\mathbf{X}[i,:]\|_{2}$, for $i=1,\ldots,d_{\text{in}}$, to scale the corresponding column of $\mathbf{W}$, and then performs magnitude based pruning on this scaled matrix to obtain the pruning mask for the weight $\mathbf{W}$. It can be viewed as approximating $\mathbf{C}^{\frac{1}{2}}$ only by its diagonal in Equation 3. Therefore, the cross-correlation information of the input activations $\mathbf{X}$ between each dimension is completely discarded.
Similar to Wanda, in the LLM quantization literature, the state-of-the-art Activation-aware Weight Quantization (AWQ) method (lin2024awq) computes the scaled $\ell_{1}$-norm of each row of $\mathbf{X}$, i.e., $\|\mathbf{X}[i,:]\|_{1}/n$, for $i=1,\ldots,d_{\text{in}}$, to scale the corresponding column of $\mathbf{W}$, and performs quantization on the scaled matrix.
@@
The Optimal Brain Compression (OBC) method (10.5555/3600270.3600593) simplifies the original Optimal Brain Surgeon (OBS) (lecun1989optimal; hassibi1993optimal) by breaking the compression task into layer-wise sub-problems. For each layer, they use a greedy solver that sequentially prunes (or quantizes) weights. To prune each row of $\mathbf{W}$, they iteratively zero out the entry that results in the smallest incremental increase to approximation loss. However, their follow-up work (frantar2023sparsegpt) mentions that this method is hard to scale to models with billions of parameters, and further propose several approximations (e.g., prune the weights from left to right and only recalculate approximation loss impact for the weights to the right) for speedup, leading to the SparseGPT method (frantar2023sparsegpt) for pruning, and the GPTQ (frantar2022gptq) method for quantization. Similarly, GPFQ (10.5555/3546258.3546414; doi:10.1137/22M1511709) uses a greedy path-following mechanism to quantize and/or prune weights from left to right, but has theoretical guarantees.
We refer interested readers to zhu-etal-2024-survey-model; survey_quantizatoin_book; Nagel2021AWP; kim2023full; wan2023efficient for a comprehensive overview of weight quantization and pruning methods. Another line of work on structured pruning, e.g., ma2023llm; 150fac892d974a75910ecbaf36c65cd0, removes entire structured components of a network, but usually involves post-training or fine-tuning to recover the performance.
@@ 3 Motivation and Method
Many existing sparse approximation and compressive sensing algorithms can be used to solve this problem, such as Matching Pursuit (1993ITSP...41.3397M), Orthogonal Matching Pursuit (OMP) (5895106), Iterative Hard Thresholding (IHT) (blumensath2009iterative), CoSaMP (tropp2008cosamp), Basis Pursuit (chen2001atomic), Sparse Bayesian Learning methods (tipping2001sparse; wipf2004sparse). In fact, the OBC method can be viewed as the OMP method in reverse order. However, it is well-known in Compressive Sensing that such methods are too greedy and often outperformed by IHT and $\ell_{1}$ type methods.
We adopt the IHT method, since it is efficient to run and also has recovery guarantees (blumensath2009iterative). IHT is a special form of the projected gradient descent (PGD) method. Each iteration of PGD involves calculating the gradient [^3] Using the fact that $\mathbf{y}=(\mathbf{W}[i,:]\mathbf{C}^{\frac{1}{2}})^{\top}$ can simplify the gradient computation without doing SVD of $\mathbf{C}$, as explained later. of $f(\boldsymbol{\theta})$, i.e., $\nabla f(\boldsymbol{\theta}^{(t)})$, and projecting the updated weights $(\boldsymbol{\theta}^{(t)}-\eta\nabla f(\boldsymbol{\theta}^{(t)}))$ onto the constraint set. For IHT, the projection onto the constraint $\|\boldsymbol{\theta}\|_{0}\leq k$ is simply hard-thresholding, i.e., keeping the $k$ largest-magnitude elements of ($\boldsymbol{\theta}^{(t)}-\eta\nabla f(\boldsymbol{\theta}^{(t)})$, and setting the remaining elements to 0. Actually all the terms in (4) can be solved independently and in parallel. This can also be viewed as minimizing the following objective, with the constraint of (5) that each row of $\hat{\mathbf{W}}$ is $k$-sparse:
@@
[…]
@@
  Repeat   $\mathbf{Z}^{(t)}=\mathbf{\Theta}^{(t)}+\eta(\mathbf{W}-\mathbf{\Theta}^{(t)})\mathbf{C}$;   $\mathbf{\Theta}^{(t+1)}=\mathsf{Proj}_{\mathcal{C}}(\mathbf{Z}^{(t)})$;
@@ 4 Experiments
We compare the proposed AWP method with state-of-the-art methods on pruning, quantization, as well as joint pruning and quantization. As in AWQ and Wanda, we evaluate the perplexity of the compressed model on the held-out WikiText-2 (merity2016pointer) validation set.
@@ 4.1 Pruning
To compare with Wanda, SparseGPT, and Magnitude-based pruning, we follow the exact setup of the experiments in Wanda [^4] https://github.com/locuslab/wanda. More specifically, we tested on the Llama-2 7B and 13B models. The calibration data $\mathbf{X}$ is 128 sequences (each has 2048 tokens) sampled from the C4 training set (raffel2020exploring). We test pruning ratios from $\{50\%,60\%,70\%,80\%,90\%\}$. Note that larger pruning ratio corresponds to higher compression rate. For AWP, the iteration stops when the Frobenius norm of the gradient normalized by the Frobenius norm of the original weight is less than 0.0001, or 200 iterations is reached. The step size $\eta$ is set as $2/\|\mathbf{C}\|_{\mathrm{F}}$. As problem (3) is nonconvex, we initialize $\mathbf{\Theta}^{(0)}$ as the solution of Wanda, since a good initial point helps nonconvex optimization.
Table 1 shows the perplexity of the pruned Llama-2-7B model for different methods and pruning ratios, and Table 2 shows the corresponding results for the Llama-2-13B model. The results of SparseGPT and Magnitude-based pruning are from sun2023wanda. As a reference, the perplexity of the original Llama-2-7B dense model (i.e., pruning ratio = 0%) is 5.12 and the perplexity of the original Llama-2-13B dense model is 4.57.
@@
Magnitude $14.89$ $4e3$ - NaN -
SparseGPT $6.51$ $9.58$ - $1e2$ -
Wanda $6.48$ $10.09$ $70.04$ $4e3$ $1e4$
@@
Magnitude $6.37$ $11.23$ - $5e4$ -
SparseGPT $5.63$ $7.80$ - $1e2$ -
Wanda $5.59$ $7.97$ $43.06$ $1e3$ $2e4$
@@ 4.2 Quantization
We compare with state-of-the-art layer-wise weight quantization methods AWQ [^5] https://github.com/mit-han-lab/llm-awq and GPTQ [^6] https://github.com/AutoGPTQ/AutoGPTQ/tree/main. We follow the experiment setups of AWQ, which focus on weight-only grouped quantization with a group size of 128. As in AWQ, we use a small calibration set from the Pile dataset (gao2020pile) in order not to overfit to a specific downstream domain. Besides INT4 and INT3 quantization, we additionally experiment with INT2 quantization.
@@ 4.3 Joint Pruning and Quantization
An interesting finding is that, compared with INT2 quantization in Table 3, INT4 quantization with 75% pruning ratio achieves significantly better perplexity. Note that 4 bits combined with 75% pruning rate is roughly equivalent to 2 bits, as we need 1 bit to store the pruning mask. As recent efforts try to push the quantization to extremely low bits, e.g., 1.58 bit (ma2024era; wang20241), our results show that combining quantization with pruning may achieve much better performance.
@@ 5 Conclusion and Discussion
Our future directions include extending pruning to structured sparsity, e.g., NVIDIA’s 2:4 sparsity (mishra2021accelerating). Another direction is to provide theoretical guarantees of the proposed method for quantization.
@@ A.1 Convergence under Restricted Isometry Property (RIP)
We first recall the theoretical guarantees of IHT algorithm (blumensath2009iterative), which uses the iteration:
@@
[blumensath2009iterative] Given a noisy observation $\mathbf{y}=\mathbf{A}\boldsymbol{\theta}_{k}+\mathbf{e}$, where $\boldsymbol{\theta}_{k}$ is $k$-sparse. If $\mathbf{A}$ has the restricted isometry property with $\beta_{3k}<1/8$, then, at iteration $t$, IHT will recover an approximation $\boldsymbol{\theta}^{(t)}$ satisfying
@@
The definition of $\beta_{3k}$ in RIP can be found in Equation (6) of blumensath2009iterative.
@@ A.2 Convergence under Restricted Strong Convexity (RSC) and Restricted Smoothness (RSM) Property
On the other hand, instead of requiring the restricted isometry property (RIP) of $\mathbf{C}^{\frac{1}{2}}$, which is not easy to verify (wang2016average), 10.5555/2968826.2968903 provided the recovery guarantee of IHT method based on Restricted Strong Convexity (RSC) and Restricted Smoothness (RSM) properties defined below (10.5555/2968826.2968903; 10.1093/imaiai/iaz027):
@@
Based on the above two properties,  10.1093/imaiai/iaz027 states that for an objective function $f$ satisfying $(\alpha,k)$-RSC and $(\beta,k)$-RSM, IHT with step size $\eta\propto 1/\beta$, i.e., ${\boldsymbol{\theta}}^{(t+1)}=H_{k}({\boldsymbol{\theta}}^{(t)}-\eta\nabla_{\boldsymbol{\theta}}f({\boldsymbol{\theta}}^{(t)}))$, satisfies
@@
[…]